Calculate the solubility of silver chloride in a solution that is 0.110 M in NH3. Express your answer using two significant figures.
Answer #1

AgCl(s) <======> Ag+(aq) + Cl-(aq)   K_{sp} = \frac{[Ag^{+}][Cl^{-}]}{[AgCl]}=1.77 \times 10^{-10}

Ag+(aq) + 2NH3(aq) <=====> Ag(NH3)2+(aq)   K_{f} = \frac{[Ag(NH_{3})_{2}^{+}]}{[Ag^{+}][NH_{3}]^{2}}=1.7 \times 10^{7}

---------------------------------------------------------------------------------------

AgCl(s) + 2NH3(aq) <======> [Ag(NH3)2]+(aq) + Cl-

I ---- 0.110 M 0 0 Using ICE table

C ----- -2x +x +x let us suppose 'x' moles [Ag(NH3)2]+ formed

E ---- 0.110 - 2x +x +x from x moles AgCl. Since AgCl is a solid we neglect moles of AgCl.

K_{eq} = \frac{[Ag(NH_{3})_{2}^{+}][Cl^{-}]}{[AgCl][NH_{3}]^{2}}

K_{eq} = K_{f} \times K_{sp}= \frac{[Ag(NH_{3})_{2}^{+}][Cl^{-}]}{[AgCl][NH_{3}]^{2}}

K_{eq} = (1.77 \times10^{-10})\times(1.7\times 10^{7}) = \frac{[Ag(NH_{3})_{2}^{+}][Cl^{-}]}{[AgCl][NH_{3}]^{2}}

3.01 \times10^{-3} = \frac{[Ag(NH_{3})_{2}^{+}][Cl^{-}]}{[AgCl][NH_{3}]^{2}}

3.01 \times10^{-3} = \frac{[x][x]}{[0.110-2x]^{2}}

3.01 \times10^{-3} = \frac{[x]^{2}}{[0.110-2x]^{2}}

\sqrt{3.01 \times10^{-3}} = \frac{[x]}{[0.110-2x]}

0.055 = \frac{[x]}{[0.110-2x]}

0.00605 - 0.11x = x

0.00605 =1.11 x

0.0054 = x

Concentration of Ag(NH3)2+ formed =x= 0.0054 M = Concentration of AgCl soluble = x = 5.4x10-3 M